# Non-abelian Hodge correspondence over singular Kähler spaces

### v1

@@ 3.2.2. Mean value inequality on orbifolds
We have the following standard lemma based on the Stokes formula for orbifolds (see e.g. [8, Theorem 3.4.2]) and the fact that $d\omega_{\mathrm{orb}}^{n-1}=0$.
@@
The following lemma can be concluded by running the argument of [12, Appendix] in the context of orbifolds. The only difference is that the constant $\widehat{C}_{1}$ doesn’t depend on $d\omega_{\mathrm{orb}}^{n-1}$ since we assume that $\omega_{\mathrm{orb}}$ is balanced (see [12, (A.27)].
@@
Let $(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})$ be an $n$-dimensional compact balanced orbifold. Suppose that $\phi(x,t)\in C^{\infty}(Y_{\mathrm{orb}}\times[0,\infty))$ is a nonnegative function satisfying the following inequality
(3.20) $(2\sqrt{-1}\Lambda_{\omega_{\mathrm{orb}}}\partial\overline{\partial}-\frac{\partial}{\partial t})\phi\geq-\breve{C}\phi$
with $\phi(0)\in L^{1}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})$, $\phi(t)\in L^{\infty}(X\setminus\Sigma,\omega)$ and $\|\phi(t)\|_{L^{1}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})}\leq\|\phi(t_{0})\|_{L^{1}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})}$ for any $t>t_{0}\geq 0$. Then we have the following statement.
@@
(3.21) $\|\phi(t_{0}+s)\|_{L^{\infty}(Y_{\mathrm{orb}})}\leq\hat{C}_{1}(\frac{1}{s^{2n+1}}+1)\|\phi(t_{0})\|_{L^{1}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})},$
where $\hat{C}_{1}$ is a positive constant depending only on $n$, $\breve{C}$ and the uniform Sobolev constant $C_{S}(\omega_{\mathrm{orb}})$.
(2) Let $Z_{\mathrm{orb}}\subset Y_{\mathrm{orb}}$ be a closed subset and $\bar{\Omega}_{\mathrm{orb}}$ be a compact subset of $Y_{\mathrm{orb}}\setminus\tilde{\Sigma}_{\mathrm{orb}}$ and $\hat{\omega}_{\mathrm{orb}}$ be an another orbifold Hermitian metric on $Y_{\mathrm{orb}}$. If there exist positive constants $\delta_{0}$ and $C_{\bar{\Omega}_{|}orb}$ such that $\overline{B_{\delta_{0}}(\bar{\Omega})}\subset Y_{\mathrm{orb}}\setminus\Sigma_{\mathrm{orb}}$ and $C_{\bar{\Omega}_{\mathrm{orb}}}^{-1}\hat{\omega}_{\mathrm{orb}}\leq\omega_{\mathrm{orb}}\leq C_{\bar{\Omega}_{\mathrm{orb}}}\hat{\omega}_{\mathrm{orb}}$ on $\overline{B_{\delta_{0}}(\bar{\Omega}_{\mathrm{orb}})}$, where $\overline{B_{\delta}(\bar{\Omega}_{\mathrm{orb}})}$ is the $\delta$-neighborhood of $\bar{\Omega}_{\mathrm{orb}}$ with respect to $\hat{\omega}_{\mathrm{orb}}$. Then,
(3.22) $\sup_{\overline{B_{\delta_{1}}(\bar{\Omega}_{\mathrm{orb}})}\times[0,T]}\phi(x,t)\leq\hat{C}_{2}((\delta_{2}-\delta_{1})^{-1}+1)^{2n+1}(\int_{0}^{T}\|\phi(t)\|_{L^{1}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})}dt+\|\phi(0)\|_{L^{\infty}(\overline{B_{\delta_{2}}(\bar{\Omega}_{\mathrm{orb}})})})$
for all $0\leq\delta_{1}<\delta_{2}\leq\delta_{0}$, where $\hat{C}_{2}$ is a constant depending only on $n$, $\breve{C}$, the Sobolev constant $C_{S}(\omega_{\mathrm{orb}})$ and $\operatorname{Vol}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})$.
@@
The following estimates are proved by Simpson for manifolds and the proof is same by restriction to the orbifold chart $U_{i}$ ([97, Lemma 6.1]).
@@
for any $0\leq t_{1}\leq t_{2}$, where $\hat{C}_{1}$ is a constant independent of $\epsilon$. Since $\omega_{\mathrm{orb},\epsilon,k}$ is strictly positive outside $f_{\mathrm{orb}}^{-1}(\Sigma)$ and so does $\omega_{\mathrm{orb},\epsilon}$. Building on the uniform Sobolev inequalities (3.18), (3.17), (3.26) Lemma 3.12 and Lemma 3.13, we have
@@
There exists constant $\tilde{C}_{1}$ which independent of $\epsilon$, such that
(3.27) $\||\Phi(H_{\mathrm{orb},\epsilon}(t_{0}+s),\omega_{\mathrm{orb},\epsilon})|_{H_{\epsilon}(t_{0}+s)}\|_{L^{\infty}(Y_{\mathrm{orb}})}\leq\tilde{C}_{1}(\frac{1}{s^{2n+1}}+1)\||\Phi(H_{\mathrm{orb},\epsilon}(t_{0}),\omega_{\mathrm{orb},\epsilon})|_{H_{\mathrm{orb},\epsilon}(t_{0})}\|_{L^{1}(Y_{\mathrm{orb}})}$
for any $s>0$ and $t_{0}\geq 0$. Let $\bar{\Omega}_{\mathrm{orb}}$ be a compact subset of $Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma)$, there exists a constant $\tilde{C}_{2}(\bar{\Omega}_{\mathrm{orb}})$ which independent of $\epsilon$, such that
(3.28) $|\Phi(H_{\mathrm{orb},\epsilon}(t),\omega_{\mathrm{orb},\epsilon})|_{H_{\mathrm{orb},\epsilon}(t)}(x)\leq\tilde{C}_{2}(\bar{\Omega}_{\mathrm{orb}})$
for any $(x,t)\in\bar{\Omega}_{\mathrm{orb}}\times[0,\infty)$.
@@
(3.30) $\log(\frac{1}{2r}(\operatorname{tr}h_{\mathrm{orb},\epsilon}(t)+\operatorname{tr}h_{\mathrm{orb},\epsilon}(t)^{-1}))\leq|s_{\epsilon}(t)|_{\hat{H}}\leq r^{\frac{1}{2}}\log(\operatorname{tr}h_{\mathrm{orb},\epsilon}(t)+\operatorname{tr}h_{\mathrm{orb},\epsilon}(t)^{-1}),$
@@
(3.32) $\log(\operatorname{tr}h_{\mathrm{orb},\epsilon}(t)+\operatorname{tr}h_{\mathrm{orb},\epsilon}^{-1}(t))-\log(2r)\leq 2\tilde{C}_{2}(\bar{\Omega}_{\mathrm{orb}})t$
@@
By using the uniform local $C^{0}$-estimate (3.33) and following the argument in [73, Lemma 2.5], we can derive uniform local $C^{\infty}$-estimates of $h_{\mathrm{orb},\epsilon}(t)$. The only difference is the Riemann curvature replacing by the Chern curvature of $\omega_{\mathrm{orb},\epsilon}$, so we omit the proof.
@@
Let $\bar{\Omega}_{\mathrm{orb}}$ be a compact subset of $Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma)$ and $T>0$. Then for every integer $j\geq 0$, there exists a constant $\bar{C}_{j+2}(\bar{\Omega}_{\mathrm{orb}},T)$ which independent of $\epsilon$, such that
(3.34) $\max_{(x,t)\in\bar{\Omega}_{\mathrm{orb}}\times[0,T]}|\nabla_{\hat{H}_{\mathrm{orb}}}^{j+2}h_{\mathrm{orb},\epsilon}(t)|_{\hat{H}_{\mathrm{orb}},\omega_{\mathrm{orb},\epsilon}}(x)\leq\hat{C}_{j+2}$
@@
Note that $\omega_{\mathrm{orb},\epsilon}$ converges to $f_{\mathrm{orb}}^{*}\omega$ from the construction (3.5) in $C_{loc}^{\infty}(Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma))$ and
$\lambda=\lim\limits_{\epsilon\rightarrow 0}\lambda_{\epsilon}=\frac{2\pi}{\int_{Y_{\mathrm{orb}}}(f_{\mathrm{orb}}^{*}\omega)^{n}}\mu_{(f^{*}\omega_{0},\cdots,f^{*}\omega_{n-2})}(E_{\mathrm{orb}})=\frac{2\pi}{\int_{Y_{\mathrm{orb}}}(f_{\mathrm{orb}}^{*}\omega)^{n}}\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{X})$
where the last equality follows from (3.2) and $\hat{H}$ denotes the descend of $\hat{H}_{\mathrm{orb}}|_{Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma)}$. From the above local uniform $C^{\infty}$-bounds on $H_{\mathrm{orb},\epsilon}(t)$, we get that, by choosing a subsequence, $H_{\mathrm{orb},\epsilon}(t)$ converges to smooth metrics $H_{\mathrm{orb}}(t)$ in $C_{loc}^{\infty}(Y_{\mathrm{orb}},f_{\mathrm{orb}}^{-1}(\Sigma))$ as $\epsilon\rightarrow 0$, which satisfies (3.9). Since $f_{\mu}$ is étale outside $f_{\mu}^{-1}(\Sigma)$, by descending $H_{\mathrm{orb}}(t)|_{Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma)}$ to a smooth Hermitian metric $H(t)$ of $\mathcal{E}_{X_{\mathrm{reg}}}|_{X\setminus\Sigma}$, the proof of Proposition 3.8 is complete.
@@ 3.3. Limiting behavior of $H(t)$
Note that $(X\setminus\Sigma,\omega)$ is a non-compact balanced manifold since $\omega^{n-1}=\omega_{0}\wedge\cdots\wedge\omega_{n-2}$. Then $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})|_{X\setminus\Sigma}$ can be seen as a Higgs bundle on $(X\setminus\Sigma,\omega)$. For simplicity, we simply write $\theta_{X_{\mathrm{reg}}}$ as $\theta$ at the moment. Let us recall some basic lemmas.
Lemma 3.17 (Cutoff functions, see e.g. [91, Lemma 5.6]).
Let $(X,\omega)$ be a compact Hermitian space of dimension $n$ and $\Sigma\subset X$ be an analytic subspace of codimension $\geq k$. Then there exists cutoff functions $\{\rho_{i}\}_{i\in N}$ ($supp(\rho_{i})\subset X\setminus\Sigma$) increasing to $1_{X\setminus\Sigma}$ and a positive constant $C$ such that $\int_{X}|d\rho_{i}|^{2k}_{\omega}+|\partial\bar{\partial}\rho_{i}|^{k}_{\omega}\omega^{n}\leq C$ for all $i\in N$ and
(3.35) $\lim_{i\rightarrow+\infty}\int_{X}|d\rho_{i}|^{2k-\delta}_{\omega}+|\partial\bar{\partial}\rho_{i}|^{k-\delta}_{\omega}\omega^{n}=0$
for any $\delta>0$ small.
@@
In lightning of Proposition 3.18 and Lemma 3.17, the uniform Sobolev inequality holds for $(X\setminus\Sigma,\omega)$.
Using the above cutoff functions and Lemma 3.18, we obtain the following mean value inequality by applying the standard Moser’s iteration as in [73, Lemma 5.2] and the stokes formula.
Lemma 3.19 (Mean value formula).
Let $(X,\omega)$ be a compact balanced space of dimension $n$ and $\Sigma\subset X$ be an analytic subspace of codimension $\geq 2$. Suppose that $u\in C^{2}(X\setminus\Sigma)$ is a bounded positive function satisfying $\Delta_{\omega}u\geq-A_{0}$. Then there exists a positive constant $\widetilde{C}$ depends only on $C_{S}$, $A_{0}$ and $\operatorname{Vol}(X,\omega)=\int_{X_{\mathrm{reg}}}\frac{\omega^{n}}{n!}$ such that
@@
These lemmas are used to take the roles of three assumptions in [97] and then we can follow a strategy of [97] as in [73, Section 4]. The remaining of the proof is a slight modification of [73, Section 4] to our context. We outline the argument for the completeness.
3.3.1. The Donaldson’s functional
Let $K$ and $H$ are two Hermitian metrics on $\mathcal{E}|_{X\setminus\Sigma}$, and set $\exp{s}=h=K^{-1}H$. By Lemma 3.1 in [97], we have

### v2

@@ 3.2.2. Mean value inequality on orbifolds
We have the following standard result based on the Stokes formula for orbifolds (see e.g. [8, Theorem 3.4.2]) and the fact that $d\omega_{\mathrm{orb}}^{n-1}=0$.
@@
The proof of the following lemma proceeds by directly adapting the argument in [12, Appendix] to the orbifold setting. The only difference is that the constant $\widehat{C}_{1}$ does not depend on $d\omega_{\mathrm{orb}}^{n-1}$ since we assume that $\omega_{\mathrm{orb}}$ is balanced (see [12, (A.27)].
@@
Let $(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})$ be an $n$-dimensional compact balanced orbifold. Suppose that $\phi_{\mathrm{orb}}(x,t)\in C^{\infty}(Y_{\mathrm{orb}}\times[0,\infty))$ is a nonnegative orbifold function satisfying the following inequality
(3.20) $(2\sqrt{-1}\Lambda_{\omega_{\mathrm{orb}}}\partial\overline{\partial}-\frac{\partial}{\partial t})\phi_{\mathrm{orb}}\geq-\breve{C}\phi_{\mathrm{orb}}$
with $\phi_{\mathrm{orb}}(0)\in L^{1}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})$ and $\|\phi_{\mathrm{orb}}(t)\|_{L^{1}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})}\leq\|\phi_{\mathrm{orb}}(t_{0})\|_{L^{1}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})}$ for any $t>t_{0}\geq 0$. Then we have the following statement:
@@
(3.21) $\|\phi_{\mathrm{orb}}(t_{0}+s)\|_{L^{\infty}(Y_{\mathrm{orb}})}\leq\breve{C}_{1}(\frac{1}{s^{2n+1}}+1)\|\phi_{\mathrm{orb}}(t_{0})\|_{L^{1}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})},$
where $\breve{C}_{1}$ is a positive constant depending only on $n$, $\breve{C}$ and the uniform Sobolev constant $C_{S}$.
(2) Let $Z_{\mathrm{orb}}\subset Y_{\mathrm{orb}}$ be a closed subset, $\bar{\Omega}_{\mathrm{orb}}$ be a compact subset of $Y_{\mathrm{orb}}\setminus Z_{\mathrm{orb}}$, and $\hat{\omega}_{\mathrm{orb}}$ be another orbifold Hermitian metric on $Y_{\mathrm{orb}}$. If there exist positive constants $\delta_{0}$ and $C_{\bar{\Omega}_{\mathrm{orb}}}$ such that $\overline{B_{\delta_{0}}(\bar{\Omega}_{\mathrm{orb}})}\subset Y_{\mathrm{orb}}\setminus Z_{\mathrm{orb}}$ and $C_{\bar{\Omega}_{\mathrm{orb}}}^{-1}\hat{\omega}_{\mathrm{orb}}\leq\omega_{\mathrm{orb}}\leq C_{\bar{\Omega}_{\mathrm{orb}}}\hat{\omega}_{\mathrm{orb}}$ on $B_{\delta_{0}}(\bar{\Omega}_{\mathrm{orb}})$, where $B_{\delta}(\bar{\Omega}_{\mathrm{orb}})$ is the $\delta$-neighborhood of $\bar{\Omega}_{\mathrm{orb}}$ with respect to $\hat{\omega}_{\mathrm{orb}}$. Then for all $0\leq\delta_{1}<\delta_{2}\leq\delta_{0}$,
(3.22) $\sup_{\overline{B_{\delta_{1}}(\bar{\Omega}_{\mathrm{orb}})}\times[0,T]}\phi_{\mathrm{orb}}(x,t)\leq\breve{C}_{2}((\delta_{2}-\delta_{1})^{-1}+1)^{2n+1}(\int_{0}^{T}\|\phi_{\mathrm{orb}}(t)\|_{L^{1}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})}dt+\|\phi_{\mathrm{orb}}(0)\|_{L^{\infty}(B_{\delta_{2}}(\bar{\Omega}_{\mathrm{orb}}))}),$
where $\breve{C}_{2}$ is a constant depending only on $n$, $\breve{C}$, the Sobolev constant $C_{S}$ and $\operatorname{Vol}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})$.
@@
The following estimates were proved by Simpson for manifolds, and the proof extends to the orbifold setting by working locally on each orbifold chart $U_{\alpha}$ (cf. [100, Lemma 6.1]).
@@
for any $0\leq t_{1}\leq t_{2}$, where $\hat{C}_{1}$ is a constant independent of $\epsilon$. Since $\omega_{\mathrm{orb},k,\epsilon}$ is strictly positive outside $f_{\mathrm{orb}}^{-1}(\Sigma)$, the same holds for $\omega_{\mathrm{orb},\epsilon}$. By combining the uniform Sobolev inequalities (3.18), (3.17) with (3.26), and applying Lemma 3.13 and Lemma 3.14, we have
@@
There exists a constant $\grave{C}_{1}$ which is independent of $\epsilon$, such that
(3.27) $\|\Phi(H_{\mathrm{orb},\epsilon}(t_{0}+s),\omega_{\mathrm{orb},\epsilon})\|_{L^{\infty}(Y_{\mathrm{orb}},H_{\mathrm{orb},\epsilon}(t_{0}+s))}\leq\grave{C}_{1}(\frac{1}{s^{2n+1}}+1)\|\Phi(H_{\mathrm{orb},\epsilon}(t_{0}),\omega_{\mathrm{orb},\epsilon})\|_{L^{1}(Y_{\mathrm{orb}},H_{\mathrm{orb},\epsilon}(t_{0}))}$
for any $s>0$ and $t_{0}\geq 0$. Furthermore, for any compact subset $\bar{\Omega}_{\mathrm{orb}}$ of $Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma)$, there exists a constant $\grave{C}_{2}(\bar{\Omega}_{\mathrm{orb}})$ independent of $\epsilon$, such that
(3.28) $|\Phi(H_{\mathrm{orb},\epsilon}(t),\omega_{\mathrm{orb},\epsilon})|_{H_{\mathrm{orb},\epsilon}(t)}(x)\leq\grave{C}_{2}(\bar{\Omega}_{\mathrm{orb}})$
for all $(x,t)\in\bar{\Omega}_{\mathrm{orb}}\times[0,\infty)$.
@@
(3.30) $\log(\frac{1}{2r}(\operatorname{tr}h_{\mathrm{orb},\epsilon}(t)+\operatorname{tr}h_{\mathrm{orb},\epsilon}(t)^{-1}))\leq|s_{\mathrm{orb},\epsilon}(t)|_{\hat{H}}\leq r^{\frac{1}{2}}\log(\operatorname{tr}h_{\mathrm{orb},\epsilon}(t)+\operatorname{tr}h_{\mathrm{orb},\epsilon}(t)^{-1}),$
@@
(3.32) $\log(\operatorname{tr}h_{\mathrm{orb},\epsilon}(t)+\operatorname{tr}h_{\mathrm{orb},\epsilon}^{-1}(t))-\log(2r)\leq 2\grave{C}_{2}(\bar{\Omega}_{\mathrm{orb}})t$
@@
By using the uniform local $C^{0}$-estimate (3.33) and following the argument in [74, Lemma 2.5], we can derive uniform local $C^{\infty}$-estimates of $h_{\mathrm{orb},\epsilon}(t)$. The only difference is that the Riemann curvature is replaced by the Chern curvature of $\omega_{\mathrm{orb},\epsilon}$, so we omit the proof.
@@
Let $\bar{\Omega}_{\mathrm{orb}}$ be a compact subset of $Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma)$ and $T>0$. Then for every integer $j\geq 0$, there exists a constant $\bar{C}_{j+2}=\bar{C}_{j+2}(\bar{\Omega}_{\mathrm{orb}},T)$ which is independent of $\epsilon$, such that
(3.34) $\max_{(x,t)\in\bar{\Omega}_{\mathrm{orb}}\times[0,T]}|\nabla_{\hat{H}_{\mathrm{orb}}}^{j+2}h_{\mathrm{orb},\epsilon}(t)|_{(\hat{H}_{\mathrm{orb}},\omega_{\mathrm{orb},\epsilon})}(x)\leq\bar{C}_{j+2}$
@@
Note that $\omega_{\mathrm{orb},\epsilon}$ converges to $f_{\mathrm{orb}}^{*}\omega$ by the construction (3.5) in $C_{loc}^{\infty}(Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma))$ and
$\lambda=\lim\limits_{\epsilon\rightarrow 0}\lambda_{\epsilon}=\frac{2\pi}{\int_{Y_{\mathrm{orb}}}(f_{\mathrm{orb}}^{*}\omega)^{n}}\mu_{(f^{*}\omega_{0},\cdots,f^{*}\omega_{n-2})}(E_{\mathrm{orb}})=\frac{2\pi}{\int_{Y_{\mathrm{orb}}}(f_{\mathrm{orb}}^{*}\omega)^{n}}\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{X}),$
where the last equality follows from (3.2) and $\hat{H}$ denotes the descent of $\hat{H}_{\mathrm{orb}}|_{Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma)}$. From the above local uniform $C^{\infty}$-bounds on $H_{\mathrm{orb},\epsilon}(t)$, we get, by choosing a subsequence, that $H_{\mathrm{orb},\epsilon}(t)$ converges in $C_{loc}^{\infty}(Y_{\mathrm{orb}},f_{\mathrm{orb}}^{-1}(\Sigma))$ to smooth metric $H_{\mathrm{orb}}(t)$ as $\epsilon\rightarrow 0$, which satisfies (3.9). Since $f_{\alpha}$ is étale outside $f_{\alpha}^{-1}(\Sigma)$, by descending $H_{\mathrm{orb}}(t)|_{Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma)}$ to a smooth Hermitian metric $H(t)$ on $\mathcal{E}_{X_{\mathrm{reg}}}|_{X\setminus\Sigma}$, the proof of Proposition 3.8 is thereby completed.
@@ 3.3. Limiting behavior of $H(t)$
Note that $(X\setminus\Sigma,\omega)$ is a non-compact balanced manifold since $\omega^{n-1}=\omega_{0}\wedge\cdots\wedge\omega_{n-2}$. Then $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})|_{X\setminus\Sigma}$ can be seen as a Higgs bundle over $(X\setminus\Sigma,\omega)$. For simplicity, we write $\theta_{X_{\mathrm{reg}}}$ as $\theta$ for now. Let us recall some basic lemmas.
Lemma 3.18 (Cutoff functions, see e.g. [93, Lemma 5.6]).
Let $(X,\omega)$ be a compact Hermitian space of dimension $n$, and let $\Sigma\subset X$ be an analytic subspace of codimension $\geq k$. Then there exist cutoff functions $\{\rho_{i}\}_{i\in\mathbb{N}}$ with $supp(\rho_{i})\subset X\setminus\Sigma$, increasing to $1_{X\setminus\Sigma}$, and a positive constant $C^{\prime}$ such that $\int_{X}\left(|d\rho_{i}|^{2k}_{\omega}+|\partial\bar{\partial}\rho_{i}|^{k}_{\omega}\right)\omega^{n}\leq C^{\prime}$ for all $i\in\mathbb{N}$, and
(3.35) $\lim_{i\rightarrow+\infty}\int_{X}\left(|d\rho_{i}|^{2k-\delta}_{\omega}+|\partial\bar{\partial}\rho_{i}|^{k-\delta}_{\omega}\right)\omega^{n}=0$
for any sufficiently small $\delta>0$.
@@
In light of Proposition 3.9 and Lemma 3.18, the uniform Sobolev inequality holds for $(X\setminus\Sigma,\omega)$.
Using the above cutoff functions and Lemma 3.19, we obtain the following mean value inequality by applying the standard Moser iteration as in [74, Lemma 5.2] and the Stokes formula.
Lemma 3.20 (Mean value inequality).
Let $(X,\omega)$ be a compact balanced space of dimension $n$ and $\Sigma\subset X$ be an analytic subspace of codimension $\geq 2$. Suppose that $u\in C^{2}(X\setminus\Sigma)$ is a bounded positive function satisfying $\Delta_{\omega}u\geq-A_{0}$ for some positive constant $A_{0}$. Then there exists a positive constant $\widetilde{C}$ depending only on $C_{S}$, $A_{0}$ and $\operatorname{Vol}(X,\omega)=\int_{X_{\mathrm{reg}}}\frac{\omega^{n}}{n!}$ such that
@@
These lemmas are used to fulfill the roles of three assumptions in [100] and then we can follow the strategy of [100] as in [74, Section 4]. The remainder of the proof follows by a slight adaptation of the argument in[74, Section 4] to our setting. We sketch it here for completeness.
3.3.1. Donaldson’s functional
Let $K$ and $H$ be two Hermitian metrics on $\mathcal{E}|_{X\setminus\Sigma}$, and set $\exp(s)=h=K^{-1}H$. By Lemma 3.1 in [100], we have